OpenStudy (anonymous):

How can I find the vertical asymptote (s) of the curve y=x^2-4/x^2-3x-10

OpenStudy (amoodarya):

y=x^2-4/x^2-3x-10 find the zero of x^2-3x-10

OpenStudy (anonymous):

The equation can be written as: \[y = \frac{ x^2 - 4 }{ x^2 - 3x -10 }\]

OpenStudy (amoodarya):

x^2-3x-10=(x-5)(x+2) x=5 is a vertical asymptote but x=-2 is not vertical asymptote because the lim f(x) x-->-2 is not infinity